FOURIER TRANSFORM SPECTROSCOPY OF BaO: ANALYSIS OF A^{1}Σ^{+} − X^{1}Σ^{+} CHEMILUMINES-CENCE

\underline{HONGZHI LI}, CRISTIAN FOCSA, BERNARD PINCHEMEL, PETER F. BERNATH and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

The A^{1}Σ^{+} − X^{1}Σ^{+} emission spectrum of BaO from a Broida-type oven was revisited using a Fourier transform spectrometer. A total of 82 vibrational bands of ^{138}BaO were measured in the range of 8,900 to 21,000 cm^{−1}at a resolution of 0.004 cm^{−1}. The vibrational quantum number of the Ground state was observed up to v′′ = 20 and up to v′ = 11 for the excited A state. In addition, 72 bands from ^{137}BaO, ^{136}Bao and 135BaO isotopomers were also measured. Over 15,000 rotational lines were analyzed at a precision of about 0.005 cm^{−1}. Significantly improved spectral constants for the ground state are obtained.